# CBGBench: Fill in the Blank of Protein-Molecule Complex Binding Graph

### v2

@@ A.3 Metric Calculation
For chemical property calculation, we use functions in RDKit [60], and note that the SA is the normalized one employed in the previous studies.
@@
For MAE metrics which are related to the molecule sizes, the generative model tends to achieve higher scores when the atom number is close to that of the reference. We suppose it is reasonable, considering that the size of the molecule itself is determined by factors such as the size and shape of the pocket, and these characteristics should be captured by the models.
Appendix B Codebase Structure
@@
Appendix C Experimental Result Details
C.1 Denovo Generation
C.1.1 Substructure
@@
Functional Group Ref. LiGan 3DSBDD GraphBP Pocket2Mol TargetDiff DiffSBDD DiffBP FLAG D3FG DecompDiff
c1ccccc1 0.3920 0.3464 0.3109 0.0133 0.3794 0.2729 0.0073 0.1962 0.2223 0.3002 0.3173
NC=O 0.1465 0.0998 0.1488 0.1330 0.1098 0.1520 0.1985 0.0108 0.1880 0.1597 0.1660
O=CO 0.1192 0.1549 0.1219 0.1888 0.2906 0.3085 0.5787 0.4440 0.1736 0.2078 0.2125
c1ccncc1 0.0452 0.0546 0.0769 0.0000 0.0305 0.0427 0.0194 0.0010 0.0365 0.0323 0.0527
c1ncc2nc[nH]c2n1 0.0340 0.1028 0.0090 0.0000 0.0000 0.0001 0.0000 0.0000 0.0009 0.0114 0.0094
NS(=O)=O 0.0303 0.0127 0.0000 0.0000 0.0000 0.0000 0.0073 0.0000 0.0000 0.0000 0.0057
O=P(O)(O)O 0.0223 0.0490 0.0814 0.0000 0.0150 0.0241 0.0000 0.0000 0.0014 0.0212 0.0418
OCO 0.0191 0.0897 0.0915 0.6064 0.0964 0.0855 0.0145 0.3289 0.2471 0.0751 0.0443
c1cncnc1 0.0178 0.0197 0.0148 0.0000 0.0030 0.0069 0.0000 0.0000 0.0135 0.0410 0.0134
c1cn[nH]c1 0.0162 0.0000 0.0291 0.0000 0.0000 0.0065 0.0000 0.0000 0.0014 0.0000 0.0128
C.1.2 Interaction
@@
Table. 14 and  14 provide detailed interaction pattern analysis. Most models captured good interaction patterns, generating a substantial amount of hydrophobic and hydrogen interactions. DiffSBDD excessively generated hydrophobic interactions. Additionally, for interactions such as $\pi$-stacking, $\pi$-cation, and halogen interactions, which do not exist in the reference molecules, the models could probabilistically generate molecules capable of producing such interactions with the protein.
@@
C.1.3 Geometry
Here we give detailed bond length and bond angle distribution of the evaluated method’s generated molecules and the reference, in Table. 16 and  16. It shows that the
@@ C.1.3 Geometry
$\mathrm{JSD}_{\mathrm{BA}}$ LiGan 3DSBDD GraphBP Pocket2Mol TargetDiff DiffSBDD DiffBP FLAG D3GF DecompDiff
C#C-C 0.6704 0.4838 0.7507 0.6477 0.6845 0.6788 0.7204 0.4591 0.7027 0.8174
C-C#N 0.8151 0.2980 0.8326 0.5830 0.7437 0.7388 0.7928 0.3461 0.7120 0.7254
C-C-C 0.5260 0.2189 0.5015 0.4663 0.2955 0.3825 0.5234 0.3439 0.3703 0.2306
C-C-N 0.5102 0.2934 0.4975 0.4790 0.2738 0.4265 0.5189 0.3650 0.3592 0.1987
C-C-O 0.5198 0.3279 0.5216 0.5078 0.3335 0.3930 0.5327 0.3710 0.4021 0.2124
C-C=C 0.4657 0.2701 0.4430 0.2826 0.1815 0.3163 0.5047 0.2830 0.2706 0.2215
C-C=N 0.4441 0.4159 0.4376 0.3507 0.2075 0.3185 0.5171 0.3353 0.3205 0.2094
C-N-C 0.5209 0.3176 0.4586 0.3981 0.2915 0.4168 0.5378 0.4237 0.3597 0.1952
C-N-N 0.5889 0.2847 0.5403 0.4997 0.2626 0.4022 0.6005 0.4161 0.3505 0.2825
C-N-O 0.7019 0.4996 0.6338 0.6173 0.3263 0.4653 0.7070 0.5560 0.4943 0.3120
C-N=C 0.3646 0.4011 0.5133 0.3728 0.3105 0.3191 0.4433 0.4085 0.4517 0.3467
C-N=N 0.3597 0.6214 0.7639 0.7062 0.4400 0.4212 0.8326 0.7023 0.7380 0.3917
C-O-C 0.5111 0.4259 0.4872 0.4204 0.2865 0.3786 0.5478 0.4606 0.3765 0.1882
C-O-N 0.7893 0.3465 0.6637 0.6140 0.4312 0.5485 0.7520 0.6257 0.4988 0.4064
C=C-N 0.4580 0.4140 0.5415 0.3732 0.2359 0.3223 0.4075 0.3361 0.3226 0.2574
C=C=C 0.7593 0.6866 0.7793 0.7373 0.7445 0.7549 0.7752 0.7419 0.7603 0.7703
N#C-C 0.8151 0.2980 0.8326 0.5830 0.7437 0.7388 0.7928 0.3461 0.7120 0.7254
N-C-N 0.5157 0.3795 0.4179 0.5544 0.3058 0.4409 0.6764 0.4316 0.4464 0.2994
N-C-O 0.4713 0.2673 0.6054 0.5879 0.3926 0.4346 0.5923 0.4089 0.4987 0.3029
N-C=N 0.4598 0.3670 0.3450 0.3986 0.2175 0.3558 0.5498 0.2654 0.3531 0.2593
N-C=O 0.5275 0.3900 0.4285 0.2347 0.2664 0.3690 0.5719 0.3636 0.3695 0.1197
N-N-O 0.8326 0.6048 0.7791 0.7639 0.5862 0.5912 0.7447 0.7117 0.6875 0.4831
N=C-N 0.4598 0.3670 0.3450 0.3986 0.2175 0.3558 0.5498 0.2654 0.3531 0.2593
O=C-N 0.5275 0.3900 0.4285 0.2347 0.2664 0.3690 0.5719 0.3636 0.3695 0.1197
C.2 Lead Optimization
We here give the details of interaction analysis. Since these tasks have provided a partial of the molecules as the context, it is hard to tell the superiority of generated substructures or geometries. Table. 18 and 18 gives details. it shows that
@@ C.3 Validity
LiGAN 0.42 10
3DSBDD 0.54 9
GraphBP 0.66 8
Pocket2Mol 0.75 5
@@
DiffSBDD 0.71 6
DiffBP 0.78 3
FLAG 0.68 7
D3FG 0.77 4
DecompDiff 0.89 2
Validity is an important metric to evaluate whether the molecules generated by the models are valid. There are two methods to reconstruct the 3D positions of the atoms into a molecule with bonds, one is used in TargerDiff, Pocket2Mol and 3DSBDD, which we name as Refine; The second is to use Openbabel [61] the software, used in DiffSBDD. However, using these methods to reconstruct molecules always carries the risk of broken bonds, which makes the strategy of selecting connected atoms to form fragments and ultimately the final molecule crucial. Here, we define a chemically valid molecule as one where the number of atoms in the largest fragment is greater than 85% of the total number of atoms, and we use this as the criterion for the reconstruction. Additionally, we employed a ‘Refine + Openbabel’ strategy, where if refinement is unsuccessful, Openbabel is added as a reconstruction method. Under this definition, the validity of various methods in the de novo task is shown in Table. 19. DecompDiff and TargetDiff have the highest validity.

### v3

@@ B.3 Metric Calculation
For chemical property calculation, we use functions in RDKit (RDKit, online, ), and note that the SA is the normalized one employed in the previous studies.
@@
For MAE metrics that are related to the molecule sizes, the generative model tends to achieve higher scores when the atom number is close to that of the reference. We suppose it is reasonable, considering that the size of the molecule itself is determined by factors such as the size and shape of the pocket, and these characteristics should be captured by the models.
B.4 Discussion on LBE.
- Is it possible that smaller molecules exhibit higher efficacy, thus leading LBE to be meaningless? (Reynolds et al., 2008)
Meaning.  The introduction of Ligand Binding Efficiency (LBE) is motivated by observations in Appendix B.2, where we identified a strong positive correlation between the size of the generated molecules and the Vina docking energy. To address this size-related bias, LBE was incorporated as an additional metric. This approach is consistent with the basic principles of energy calculation, where the total binding energy is derived from the sum of interaction energies contributed by each atom, i.e. $E_{\mathrm{vina}}\sim E_{\mathrm{bind}}=\sum_{i=1}^{N_{\mathrm{atom}}}E_{i}$. Thus, using LBE as a measure of the average per-atom contribution to binding affinity is both logical and appropriate, as $\mathrm{LBE}=-\frac{E_{\mathrm{bind}}}{N_{\mathrm{atom}}}$. Furthermore, numerous studies in medicinal chemistry have highlighted Ligand Efficiency as a critical evaluation criterion, advocating for its widespread use as an effective metric for assessing molecular binding affinity (Hopkins et al., 2014b; Kenny, 2017; Murray et al., 2010).
Table 11: Ratio of generated molecule’s size and the corresponding LBE.
Method Mol Size [1,3] (3,10] (10,20] (20,30] >30
LiGAN Ratio(%) 0.01 13.61 40.09 36.77 9.59
LBE 0.37 0.42 0.44 0.35 0.30
3DSBDD Ratio(%) 0.01 15.48 54.44 22.63 7.42
LBE 0.39 0.42 0.41 0.33 0.31
GraphBP Ratio(%) 0.00 0.77 91.29 6.07 1.84
LBE - 0.36 0.33 0.21 0.20
Pocket2Mol Ratio(%) 0.08 24.22 48.83 20.19 6.64
LBE 0.43 0.45 0.44 0.33 0.30
TargetDiff Ratio(%) 0.00 10.24 28.17 39.51 22.05
LBE - 0.44 0.44 0.31 0.29
DiffSBDD Ratio(%) 0.13 21.54 41.43 28.01 8.86
LBE 0.37 0.38 0.30 0.24 0.22
DiffBP Ratio(%) 0.00 7.29 32.19 43.08 17.42
LBE - 0.41 0.42 0.31 0.29
FLAG Ratio(%) 0.04 39.01 54.29 6.52 0.13
LBE 0.40 0.35 0.33 0.24 0.19
D3FG Ratio(%) 0.00 8.31 61.70 27.72 2.55
LBE - 0.44 0.45 0.32 0.25
DecompDiff Ratio(%) 0.00 9.99 35.50 33.98 20.51
LBE - 0.38 0.38 0.33 0.30
MolCraft Ratio(%) 0.00 9.18 30.97 39.62 20.21
LBE - 0.44 0.41 0.33 0.29
VoxBind Ratio(%) 0.00 6.10 36.64 43.39 15.85
LBE - 0.43 0.41 0.34 0.26
Resonability. It can be usually concluded from Reynolds et al. (2008) that ‘smaller molecules tend to have higher LBE’, while we believe this conclusion is incorrect. We claim that different protein pockets exhibit a preference for molecule sizes. If this conclusion were true, it would imply that molecules consisting of a single atom would have the optimal LBE? To explore this issue, in Table. 11, we give statistics of molecule sizes v.s. LBE of methods evaluated. It shows that such molecules with extreme sizes (less than 4) that make the LBE meaningless are scarcely generated since they have a small ratio. Ligands with sizes less than 20 usually have better LBE because of the preference of the pockets, as a consequence of protein binding sites being limited in their size, thus generating the preference of binding molecules’ sizes. Larger ligand with sizes of more than 20 gets decreased LBE because it becomes increasingly difficult to form optimal interactions with every site on the protein without introducing unfavorable ligand strain.
Little Effects of Extreme Samples.  Besides, we use a weighted ranking with equal weights as previous benchmarks (Wang et al., 2022), to give an unbiased evaluation. Given that LBE accounts for one-twelfth of the total weight in interaction analysis, the impact of failed samples in LBE on calculating the LBE metrics and final ranking is minimal. Therefore, we claim that the inclusion of a minimal number of invalid LBE calculations will not render the overall ranking ineffective.
Table 12: Different methods’ evaluation aspects in the published papers.
Method Substructure Geometry Chem Property Interaction
LiGAN Figure. S6 Figure. S7, S8, S9 Figure. 3, S3, S4, S5 Figure. S13, S14
Pocket2Mol Table. 2 Table. 3; Figure. 4 Table. 1 Table. 1
GraphBP - Table. 2; Figure. 5 - Table. 1; Figure. 2
TargetDiff Table. 2 Table. 1; Figure. 2 Table. 3 Table. 3, Figure. 4
DiffBP Table. 3 - Table. 2 Table. 1
DiffSBDD - Figure. 8, 9 Table. 1 Table. 1, 2; Figure. 2
FLAG Table. 3 Table. 2; Figure. 4 Table. 1 Table. 1
D3FG Table. 1, 3; Figure. 3 Table. 2 Table. 4 Table. 4
DecompDiff - Table. 1, 2; Figure. 3 Table. 3 Table. 3
MolCraft Table. 1 Table. 2 Table. 2 Table. 2
VoxBind - Figure. 7 Table. 1 Table. 1; Figure. 5
B.5 Previously-Used Metrics for Evaluation
The aspects of evaluation are commonly used with different methods, while they hardly give a very comprehensive evaluation. We give a brief review of each method’s evaluation on these aspects in Table. 12.
Appendix C Codebase Structure
@@
Appendix D Experimental Implementation Details
Experiments are conducted based on Pytorch 2.0.1 on a hardware platform with Intel(R) Xeon(R) Gold 6240R @ 2.40GHz CPU and NVIDIA A100 GPU. We give the detailed parameters for training the included models in Table. 13. Note that most of the hyperparameters’ combinations are directly the officially-provided ones. Specifically, The methods included in our codebase are either self-trained or trained by us when the official repository does not provide a pretrained checkpoint. For methods that are not included, we validate them using the pretrained models that have been provided. In the Lead-Optimization tasks, the finetuning parameters are the same.
Table 13: The hyper-parameters in for training and sampling molecules for the SBDD methods. Different from GNN-based methods, CNN-based methods usually have a different hidden dimension size, so we report the highest ones. ‘Num steps’ are the sampling steps which is usually used in diffusion-based methods, including DDPMs and BFNs. For other hyper-parameters in detail, please read the config files provided in the supplementary materials from ./configs/denovo/train.
Method Batch Size Hidden Dim Layer num Learning rate Optimizer Num Steps
LiGan 8 128 4 1E-05 RMSprop /
3DSBDD 4 256 6 1E-04 Adam /
GraphBP 16 64 6 1E-04 Adam /
Pocket2Mol 8 256 6 1E-05 Adam /
TargetDiff 4 128 9 5E-04 Adam 1000
DiffSBDD 16 256 6 1E-03 Adam 500
DiffBP 16 256 6 1E-04 Adam 1000
FLAG 4 256 6 1E-04 Adam /
D3FG 16 128 6 1E-04 Adam 1000
DecompDiff 4 128 6 5E-04 Adam 1000
MolCraft 8 128 9 5E-04 Adam 1000
VoxBind 32 512 4 1E-05 AdamW 100
Appendix E Experimental Result Details
E.1 De novo Generation
E.1.1 Substructure
@@
Method c1ccccc1 NC=O O=CO c1ccncc1 c1ncc2nc[nH]c2n1 NS(=O)=O O=P(O)(O)O OCO c1cncnc1 c1cn[nH]c1
Ref. 0.3920 0.1465 0.1192 0.0452 0.0340 0.0303 0.0223 0.0191 0.0178 0.0162
LiGan 0.3464 0.0998 0.1549 0.0546 0.1028 0.0127 0.0490 0.0897 0.0197 0.0000
3DSBDD 0.3109 0.1488 0.1219 0.0769 0.0090 0.0000 0.0814 0.0915 0.0148 0.0291
GraphBP 0.0133 0.1330 0.1888 0.0000 0.0000 0.0000 0.0000 0.6064 0.0000 0.0000
Pocket2Mol 0.3794 0.1098 0.2906 0.0305 0.0000 0.0000 0.0150 0.0964 0.0030 0.0000
TargetDiff 0.2729 0.1520 0.3085 0.0427 0.0001 0.0000 0.0241 0.0855 0.0069 0.0065
DiffSBDD 0.0073 0.1985 0.5787 0.0194 0.0000 0.0073 0.0000 0.0145 0.0000 0.0000
DiffBP 0.1962 0.0108 0.4440 0.0010 0.0000 0.0000 0.0000 0.3289 0.0000 0.0000
FLAG 0.2223 0.1880 0.1736 0.0365 0.0009 0.0000 0.0014 0.2471 0.0135 0.0014
D3FG 0.3002 0.1597 0.2078 0.0323 0.0114 0.0000 0.0212 0.0751 0.0410 0.0000
DecompDiff 0.3173 0.1660 0.2125 0.0527 0.0094 0.0057 0.0418 0.0443 0.0134 0.0128
MolCraft 0.3960 0.1798 0.2361 0.0377 0.0001 0.0000 0.0362 0.0344 0.0053 0.0002
VoxBind 0.4705 0.1138 0.1663 0.0762 0.0001 0.0000 0.0191 0.0334 0.0110 0.0000
E.1.2 Interaction
@@
Table. 18 and  18 provide detailed interaction pattern analysis. Most models captured good interaction patterns, generating a substantial amount of hydrophobic and hydrogen interactions. DiffSBDD excessively generated hydrophobic interactions. Additionally, for interactions such as $\pi$-stacking, $\pi$-cation, and halogen interactions, which do not exist in the reference molecules, the models could probabilistically generate molecules capable of producing such interactions with the protein.
@@
MolCraft 0.4855 0.4771 0.0000 0.0246 0.0080 0.0045 0.0000
VoxBind 0.5943 0.3757 0.0000 0.0219 0.0007 0.2012 0.0000
E.1.3 Geometry
Here we give detailed bond length and bond angle distribution of the evaluated method’s generated molecules and the reference, in Table. 20 and  20. It shows that the
@@ E.1.3 Geometry
$\mathrm{JSD}_{\mathrm{BA}}$ LiGan 3DSBDD GraphBP Pocket2Mol TargetDiff DiffSBDD DiffBP FLAG D3GF DecompDiff MolCraft VoxBind
C#C-C 0.6704 0.4838 0.7507 0.6477 0.6845 0.6788 0.7204 0.4591 0.7027 0.8174 0.4922 0.6275
C-C#N 0.8151 0.2980 0.8326 0.5830 0.7437 0.7388 0.7928 0.3461 0.7120 0.7254 0.2252 0.7320
C-C-C 0.5260 0.2189 0.5015 0.4663 0.2955 0.3825 0.5234 0.3439 0.3703 0.2306 0.1926 0.2742
C-C-N 0.5102 0.2934 0.4975 0.4790 0.2738 0.4265 0.5189 0.3650 0.3592 0.1987 0.1097 0.2280
C-C-O 0.5198 0.3279 0.5216 0.5078 0.3335 0.3930 0.5327 0.3710 0.4021 0.2124 0.1277 0.3233
C-C=C 0.4657 0.2701 0.4430 0.2826 0.1815 0.3163 0.5047 0.2830 0.2706 0.2215 0.2267 0.1823
C-C=N 0.4441 0.4159 0.4376 0.3507 0.2075 0.3185 0.5171 0.3353 0.3205 0.2094 0.3175 0.2707
C-N-C 0.5209 0.3176 0.4586 0.3981 0.2915 0.4168 0.5378 0.4237 0.3597 0.1952 0.1475 0.1532
C-N-N 0.5889 0.2847 0.5403 0.4997 0.2626 0.4022 0.6005 0.4161 0.3505 0.2825 0.2702 0.1813
C-N-O 0.7019 0.4996 0.6338 0.6173 0.3263 0.4653 0.7070 0.5560 0.4943 0.3120 0.2002 0.3300
C-N=C 0.3646 0.4011 0.5133 0.3728 0.3105 0.3191 0.4433 0.4085 0.4517 0.3467 0.4092 0.3626
C-N=N 0.3597 0.6214 0.7639 0.7062 0.4400 0.4212 0.8326 0.7023 0.7380 0.3917 0.6954 0.7196
C-O-C 0.5111 0.4259 0.4872 0.4204 0.2865 0.3786 0.5478 0.4606 0.3765 0.1882 0.1113 0.3305
C-O-N 0.7893 0.3465 0.6637 0.6140 0.4312 0.5485 0.7520 0.6257 0.4988 0.4064 0.3336 0.4436
C=C-N 0.4580 0.4140 0.5415 0.3732 0.2359 0.3223 0.4075 0.3361 0.3226 0.2574 0.1760 0.2905
C=C=C 0.7593 0.6866 0.7793 0.7373 0.7445 0.7549 0.7752 0.7419 0.7603 0.7703 0.6388 0.8326
N#C-C 0.8151 0.2980 0.8326 0.5830 0.7437 0.7388 0.7928 0.3461 0.7120 0.7254 0.2252 0.7320
N-C-N 0.5157 0.3795 0.4179 0.5544 0.3058 0.4409 0.6764 0.4316 0.4464 0.2994 0.2604 0.2569
N-C-O 0.4713 0.2673 0.6054 0.5879 0.3926 0.4346 0.5923 0.4089 0.4987 0.3029 0.2612 0.4190
N-C=N 0.4598 0.3670 0.3450 0.3986 0.2175 0.3558 0.5498 0.2654 0.3531 0.2593 0.1366 0.1360
N-C=O 0.5275 0.3900 0.4285 0.2347 0.2664 0.3690 0.5719 0.3636 0.3695 0.1197 0.0378 0.2508
N-N-O 0.8326 0.6048 0.7791 0.7639 0.5862 0.5912 0.7447 0.7117 0.6875 0.4831 0.6708 0.5859
N=C-N 0.4598 0.3670 0.3450 0.3986 0.2175 0.3558 0.5498 0.2654 0.3531 0.2593 0.1366 0.1360
O=C-N 0.5275 0.3900 0.4285 0.2347 0.2664 0.3690 0.5719 0.3636 0.3695 0.1197 0.0378 0.2508
E.2 Lead Optimization
We here give the details of interaction analysis. Since these tasks have provided a partial of the molecules as the context, it is hard to tell the superiority of generated substructures or geometries. Table. 22 and 22 gives details. it shows that
@@ E.4 Validity
LiGAN 0.42 12
3DSBDD 0.54 11
GraphBP 0.66 10
Pocket2Mol 0.75 6
@@
DiffSBDD 0.71 7
DiffBP 0.78 4
FLAG 0.68 9
D3FG 0.77 5
DecompDiff 0.89 3
MolCraft 0.95 2
VoxBind 0.74 8
Validity is an important metric to evaluate whether the molecules generated by the models are valid. There are two methods to reconstruct the 3D positions of the atoms into a molecule with bonds, one is used in TargerDiff, Pocket2Mol and 3DSBDD, which we name as Refine; The second is to use Openbabel (Open Babel development team, ) the software, used in DiffSBDD. However, using these methods to reconstruct molecules always carries the risk of broken bonds, which makes the strategy of selecting connected atoms to form fragments and ultimately the final molecule crucial. Here, we define a chemically valid molecule as one where the number of atoms in the largest fragment is greater than 85% of the total number of atoms, and we use this as the criterion for the reconstruction. Additionally, we employed a ‘Refine + Openbabel’ strategy, where if refinement is unsuccessful, Openbabel is added as a reconstruction method. Under this definition, the validity of various methods in the de novo task is shown in Table. 23. DecompDiff and TargetDiff have the highest validity.